A block of mass m = 12 Kg is pushed at a constant speed, up an incline plane (angle of


inclination 30o


) though a distance of s= 4.5m and raised to a height h= 2m. The coefficient


of kinetic friction between the block and the incline is 0.41. Calculate the work done by


(a) Normal force


(b) the external force applied parallel to the incline.


(c) gravity


(d) the frictional force.

1
Expert's answer
2021-12-12T16:40:35-0500

Work done by the

(a) Normal force:


"W=hN=hmg\cos\theta=204\text{ J}."

(b) the external force applied parallel to the incline:


"W=sF=s\mu mg\cos\theta=188\text{ J}."

(c) Gravity:


"W=mgh=235\text{ J}."

(d) The frictional force:


"W=sf=s\mu mg\cos\theta=188\text{ J}."